If QR=6x+1,PR=14x-13 and PQ=5x-2,find PR
1
Expert's answer
2020-09-23T16:41:36-0400

"PQ+QR=PR"

"5x-2+6x+1=14x-13"

"11x-1=14x-13"

"11x-14x=1-13"

"-3x=-12"

"x=-12:(-3)"

"x=4"

Calculate the value of "PR":

"PR=14x-13=14*4-13=56-13=43"

Solution: "PR=43"